OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (1, 3) and is perpendicular to the line y = –x + 2?

OpenStudy (anonymous):

the options are y = –x – 4 y = x + 2 y = –x + 2 y = x – 4

OpenStudy (saifoo.khan):

perpendicular to that slope means slope of +1. y = mx + b 3 = 1(1) + b b = 3-2 b = 2 y = mx + b y = x + 2

OpenStudy (saifoo.khan):

3-1 = 1. OMG what happened to me! >_< lol.

OpenStudy (saifoo.khan):

So your answer is B.